OpenStudy (anonymous):

What is the equation of the line that contains the points (3, –6) and (–3, 0)? A. y = x – 3 B. y = –x – 3 C. y = –x + 3 D. y = x + 3

OpenStudy (anonymous):

ill give medal

OpenStudy (anonymous):

first of all you have to find the slope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know how to find the slope?

OpenStudy (anonymous):

you use the equation : (y2-y2)/(x2-x1)

OpenStudy (anonymous):

y2-y1*

OpenStudy (anonymous):

so its 3,3 right

OpenStudy (anonymous):

what is 3,3?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

you just substitute into the equation : (3,-6) (-3,0) y2 = 0 y1 = -6 x2 = -3 x1= 3

OpenStudy (anonymous):

so basically what would you get? (keeping in mind that the equation is y2-y1/x2-x1

OpenStudy (anonymous):

y=x+3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

nope, try it again look : it would be 0--6/ -3-3 = 6/-6 = -1 so the slope would be -x

OpenStudy (anonymous):

and the three would be positive

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so basically we have found that the slope is -x, but we dont know what the y-intercept is, so for now, our equation is y=-x+c where c is the unknown y-interecept, so you substitute into this equation using one of the points such as (-3,0) so 0 = - (-3) + c so 0 = 3 +c so c = -3

OpenStudy (anonymous):

ahh i get it now thank you

OpenStudy (anonymous):

you're welcome :)